A 50 N block is placed on a smooth inclined plane which is 5 meters long and 3 meters high. (a) What force parallel to the plane and applied to the block will keep the block in equilibrium? (b) What force is applied by the plane to the block? 


1
Expert's answer
2021-07-05T17:59:20-0400

a)


"F=W\sin{\alpha}\\F=50\frac{3}{5}=30\ N"


b)


"N=W\cos{\alpha}\\N=50\frac{4}{5}=40\ N"